Note. No class on Mónday, January 21, for observance Of MILR day. Consider a mortgage of $150,000 at an interest rate of 3.6% APR compounded monthly for 30 years. 1. What would your monthly payment (PI) be? 2. How much interest would you pay over the 30 years note? 3. If you pay an extra $600 each month, how long would it take to pay off the loan? Refer to #3, how much interest would you pay? 5. What would be your monthly payment if you wanted to pay off the mortgage in 10 years?
Answer #1

5) A= d[(1+i)^nt - 1]/i

A= 150000

i= 3.6%/12= 0.003

nt=12*10= 120

150000= d[(1+0.003)^120 - 1)]/0.003

450= d(0.432557)

d= 1040.325

Monthly Payment should be 1040.325